# Detecting Out-of-Distribution Objects through Class-Conditioned Inpainting

### v3

@@ 4.3 Vision-language Triplet Measurement
Figure 3: Triplet similarity relationships between (i) the original object, (ii) the inpainted outcome, and (iii) the predicted ID label. In-distribution (ID) samples demonstrate strong alignment across all three components, showing high semantic and visual similarity. Conversely, out-of-distribution (OOD) samples exhibit poor alignments, particularly between the original object and the predicted label or inpainted object, effectively differentiating ID from OOD.
@@
$\textit{similarity}{(ori,\ \hat{y})}$ $=\mathbf{cosine}\left(f_{v}(\mathbf{x_{ori}}),f_{t}(\hat{y})\right)$ (3)
$\textit{similarity}{(inp,\ \hat{y})}$ $=\mathbf{cosine}\left(f_{v}(\mathbf{x_{inp}}),f_{t}(\hat{y})\right)$
$\textit{similarity}{(ori,\ inp)}$ $=\mathbf{cosine}\left(f_{i}(\mathbf{x_{ori}}),f_{i}(\mathbf{x_{inp}})\right)$
Here, $\mathbf{cosine}$ means the cosine similarity between two vectors and $f_{v}$ and $f_{t}$ represent visual and textual embeddings from large vision-language contrastive models (e.g., CLIP (Radford et al., 2021)), while $f_{i}$ is derived from a visual contrastive model (e.g., SimCLRv2 (Chen et al., 2020)). Cosine similarity effectively measures the alignment between these representations. Specifically, $\textit{similarity}{(ori,\ inp)}$ emphasizes visual alignment for ID retention while $\textit{similarity}{(ori,\ \hat{y})}$ supports semantic alignment for OOD recognition. In contrast, $\textit{similarity}{(inp,\ \hat{y})}$ helps normalize variations across different labels. Figure 3 illustrates the relationships between these components for both ID and OOD cases.
@@
where $\bm{\alpha}$ and $\bm{\beta}$ are hyperparameters controlling the balance between visual-language and visual similarity. A higher $\text{S}_{\text{triplet}}$ suggests a greater likelihood of the object being ID, while lower values indicate OOD cases. In practice, a thresholding method can be applied for effective OOD detection depending on different tasks.
@@ 4.4 Refined Prompt for Near-OOD
Identifying near-OOD cases poses a significant challenge for conventional OOD detection methods, as unseen objects often share closely related semantic meanings with certain in-distribution (ID) categories, making separation difficult. However, off-the-shelf generative models inherently possess the ability to distinguish them due to the abundant pre-training data. Given a predefined set of ID labels, we can leverage knowledge graph or taxanomy to retrieve closely related classes or concepts that are OOD and explicitly instruct generative models not to inpaint such concepts as conditioning. For example, if the ID label is “horse”, we can retrieve highly related but OOD classes such as—donkey, zebra, mule, pony, and camel—and explicitly condition the inpainting model to generate a horse that remains distinct from these related categories. This “refined” prompting technique not only guides RONIN to synthesize inpainted outcomes that align better with the ID label but also strengthens its robustness against near-OOD samples.
@@ 5 Experiments
In-Distribution Method MS-COCO OpenImages FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ ) FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ ) PASCAL-VOC ODIN (Liang et al., 2018) 41.65 88.22 55.87 86.46 Energy Score (Liu et al., 2020) 29.48 90.26 24.57 91.24 Mahalanobis (Xiao et al., 2021) 63.30 83.26 45.22 87.11 KNN (Sun et al., 2022) 58.56 85.52 45.00 84.62 VOS (Du et al., 2022b) 48.15 88.75 54.63 83.65 SIREN (Du et al., 2022a) 64.70 78.68 66.69 75.12 MCM (Ming et al., 2022) 62.47 83.15 71.52 81.45 CLIPN (Wang et al., 2023) 43.09 85.45 41.74 89.31 TAG (Liu & Zach, 2024) 61.03 78.35 48.48 87.69 OLE (Ding & Pang, 2024) 54.23 86.13 49.13 88.07 GL-MCM (Miyai et al., 2025) 78.56 78.64 69.35 82.89 RONIN   (5 steps) 29.48 90.39 20.00 92.32 RONIN (10 steps) 29.07 91.10 20.20 92.80 RONIN (15 steps) 25.57 91.63 19.87 92.84 RONIN (20 steps) 25.36 92.31 18.91 93.10 BDD-100k ODIN (Liang et al., 2018) 96.51 55.18 95.56 57.87 Energy Score (Liu et al., 2020) 71.75 74.49 53.33 73.09 Mahalanobis (Xiao et al., 2021) 31.75 90.74 23.33 93.81 KNN (Sun et al., 2022) 35.87 86.58 31.11 92.75 VOS (Du et al., 2022b) 65.45 78.34 59.23 80.42 SIREN (Du et al., 2022a) 42.86 89.37 37.97 91.78 MCM (Ming et al., 2022) 95.56 55.82 92.22 57.05 CLIPN (Wang et al., 2023) 28.49 92.14 44.76 85.78 TAG (Liu & Zach, 2024) 46.43 90.85 55.30 76.67 OLE (Ding & Pang, 2024) 57.14 78.14 40.00 85.36 GL-MCM (Miyai et al., 2025) 96.83 50.82 94.44 58.88 RONIN   (5 steps) 27.94 92.59 26.67 91.23 RONIN (10 steps) 26.67 92.73 25.56 91.33 RONIN (15 steps) 26.03 92.83 24.44 91.65 RONIN (20 steps) 26.03 92.90 23.33 91.90
Figure 4: Visualizations of RONIN’s inpaintings. The ID inpaintings closely resemble the original objects, while the OOD ones appear distinctly different. This allows RONIN to effectively distinguish between ID and OOD objects.
@@ 5.1 Experimental Setups
Datasets.   We use Pascal-VOC (Everingham et al., 2010) and Berkerly DeepDrive (BDD-100k) (Yu et al., 2020) as ID data, and MS-COCO (Lin et al., 2014) and OpenImages (Kuznetsova et al., 2020) as OOD data, covering a wide range of indoor and outdoor scenes. We follow the splits from Du et al. (2022a, b), where overlapping categories in the ID and OOD datasets are removed. We select a subset of 200 images from the test set of BDD-100k and 400 each from the test sets of Pascal-VOC, MS-COCO, and OpenImages. Data processing details can be found in Appendix A.
@@
Baselines.   We consider three types of baselines: (1) Discriminative zero-shot approaches: MCM (Ming et al., 2022), CLIP-based ODIN (Liang et al., 2018), CLIP-based Energy Score (Liu et al., 2020), CLIPN (Wang et al., 2023), TAG (Liu & Zach, 2024), OLE (Ding & Pang, 2024) and GL-MCM (Miyai et al., 2025). (2) Alternative generative zero-shot approaches: We first generate synthetic data and then apply standard OOD detection methods such as Mahalanobis (Xiao et al., 2021) and KNN (Sun et al., 2022). (3) Training-based approaches: for a complete comparison, we also include two representative training-based object-level methods VOS (Du et al., 2022b) and SIREN (Du et al., 2022a), resulting in a total of eleven baselines. Implementation details can be found in Appendix B.
Implementation Details.    For RONIN, we perform class-wise inpainting with Stable Diffusion 2 Inpainting (Rombach et al., 2021) with 20 steps. Each object is masked with a center mask covering 0.9 of the height and width of its original bounding box. Object-label similarities are determined by the distance between OpenCLIP (Ilharco et al., 2021) features, and object-object similarities by the SimCLRv2 (Chen et al., 2020) features.
For the main experiments in subsection 5.2, we use detected objects obtained from Deformable-DETR (Zhu et al., 2020), as implemented by Du et al. (2022a), to enable direct comparisons with training-based baselines. However, it is important to note that RONIN is compatible with various types of detectors. In subsection 5.3, we demonstrate RONIN ’s performance using detections from the pre-trained open-vocabulary detector GroundingDINO (Liu et al., 2023a), showcasing its flexibility and compatibility. The Triplet Similarity (Equation 4) is computed with $\bm{\alpha}=2$ and $\bm{\beta}=1$.
5.2 Main Experimental Results
Table 1 presents the performance of RONIN and the baselines based on the DeformableDETR detections in four settings. RONIN achieves consistent and competitive performance across all four OOD scenarios, attaining the highest performance in three of them even with just 5 denoising steps. It even outperforms training-based methods VOS and SIREN, demonstrating the effectiveness of using off-the-shelf models to bypass the need for retraining with dedicated in-domain data. Using more denoising steps further improves the performance. In PascalVOC vs. OpenImages, it reduces FPR@95 by 23% over the best-performing baseline. Meanwhile, purely discriminative zero-shot approaches show fluctuating results. Alternative generative approaches combined with standard OOD detection show reasonable performance but still underperform RONIN. This further highlights the advantages of incorporating generative models for object-level OOD detection.
Figure 4 visualizes RONIN’s inpaintings using PascalVOC as the in-distribution dataset. The inpaintings for ID objects closely resemble the original ones, while those for OOD objects appear distinctly different. This allows RONIN to effectively differentiate between the ID and OOD objects.
@@ 5.3 Near-OOD Detection
Figure 5: RONIN performances over near-OOD samples with the refined prompting techique. By incorporating closely related species as supplementary exclusions, our approach attains inpaintings that align accurately with the predicted ID labels, thus achieving lower similarity scores for effective near-OOD detection.
To evaluate RONIN ’s ability to handle hard OOD objects, we developed a small near-OOD dataset where ID and OOD objects are both visually and semantically similar. Specifically, we curated 411 samples by selecting ID objects from PascalVOC and OOD objects from MS-COCO and OpenImages, focusing on closely related category ID-OOD pairs such as “horse vs. zebra”, “dog vs. fox”, and “cat vs. raccoon/tiger”. We obtained object predictions using the open-vocabulary object detector GroundingDINO (Liu et al., 2023a), yielding 245 ID objects and 265 OOD objects. As described in subsection 4.4, we apply a “refined” prompting technique that incorporates five pre-anticipated nearest OOD labels as exclusions, with the conditioning prompt for diffusion inpainting illustrated in Figure 5.
Table 2: Near-OOD detection performance. RONIN performs competitively in this setting.
FPR@95 ($\downarrow$) AUROC ($\uparrow$) ODIN (Liang et al., 2018) 19.62 94.87 Energy Score (Liu et al., 2020) 9.81 97.65 MCM (Ming et al., 2022) 31.32 93.65 CLIPN (Wang et al., 2023) 3.77 97.80 TAG (Liu & Zach, 2024) 29.06 94.02 OLE (Ding & Pang, 2024) 3.40 98.33 GL-MCM (Miyai et al., 2025) 44.91 85.66 RONIN (Simple prompting) 7.79 96.86 RONIN (Refined prompting) 5.66 98.03
Table 2 compares RONIN with selected baselines on the near-OOD dataset, showing that our proposed method achieves comparable performance in this challenging scenario, particularly when using the “refined” prompting technique. This improvement is further illustrated in Figure 5, where the additional information in the refined prompt enables RONIN to generate outcomes that align more closely with the predicted ID label, resulting in lower similarity scores and thus better near-OOD detection performance. The distinct qualitative and quantitative improvements between the two prompting techniques strengthen our hypothesis that pre-trained, off-the-shelf generative and multi-modal models already possess the capacity for fine-grained near-OOD detection, given sufficient information.
@@ 6 Ablation Studies
Figure 6: Performance-speed tradeoff of RONIN. Smaller image size and fewer denoising steps significantly improve speed with minimal performance loss. At full configuration, RONIN achieves the best performance with reasonable runtime, showcasing its adaptability to efficiency and performance needs. For relative runtime comparison, we incorporate the best-performance baseline, CLIPN, though we do not focus on runtime gaps.
@@
Performance-Speed Trade off.   In Figure 6, we analyze the trade-off between RONIN ’s performance and runtime by varying the number of denoising steps and the image resolutions used during the inpainting process. For relative performance comparison, we include CLIPN (Wang et al., 2023) as the best-performing baseline, though we do not compare absolute runtimes between the two methods. Reducing the number of denoising steps and downscaling image sizes significantly accelerates RONIN, with only minor performance degradation. These results highlight the flexibility of our approach and its potential suitability for low-resource compute environments, with reasonable performance drops.
@@
Table 3: RONIN with object-wise and class-wise inpainting. Class-wise inpainting attain better performance while having faster speed.
@@
FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ ) Object-wise Class-wise Object-wise Class-wise VOC - COCO 26.60 25.36 91.48 92.31 VOC - OpenImages 20.00 18.91 92.68 93.10 BDD - COCO 28.89 26.03 93.74 92.90 BDD - OpenImages 26.67 23.33 92.42 91.90
Class-wise vs. Object-wise Inpainting.    Table 3 compares the performance between “object-wise” inpainting strategy, where each object is masked and inpainted individually, and our default “class-wise” inpainting approach, where objects are processed together based on their labels. By sharing the inpainting process among objects with the same label, class-wise inpainting achieves comparable reconstruction quality while delivering better performance and greater efficiency compared to the object-wise strategy.

### v4

@@ 4.2 Vision-language Triplet Measurement
Figure 3: Triplet similarity relationships between (i) the original object, (ii) the inpainted outcome, and (iii) the predicted label. ID samples show strong alignments across all three, whereas OOD samples exhibit weak alignments, aiding effective OOD detection.
@@
[…]
$\textit{similarity}{(ori,\ inp)}$ $=\mathbf{cosine}\left(f_{i}(\mathbf{x_{ori}}),f_{i}(\mathbf{x_{inp}})\right)\$
Here, $\mathbf{cosine}$ means the cosine similarity between two vectors and $f_{v}$ and $f_{t}$ represent visual and textual embeddings from large vision-language contrastive models (e.g., CLIP [radford2021clip]), while $f_{i}$ is derived from a visual contrastive model (e.g., SimCLRv2 [chen2020simclrv2]). Cosine similarity effectively measures the alignment between these representations. Specifically, $\textit{similarity}{(ori,\ inp)}$ emphasizes visual alignment for ID retention while $\textit{similarity}{(ori,\ \hat{y})}$ supports semantic alignment for OOD recognition. In contrast, $\textit{similarity}{(inp,\ \hat{y})}$ helps normalize variations across different labels. Fig. 3 illustrates the relationships between these components for both ID and OOD cases.
@@
where $\bm{\alpha}$ and $\bm{\beta}$ are hyperparameters controlling the balance between visual-language and visual similarities. A higher $\text{S}_{\text{triplet}}$ suggests a greater likelihood of the object being ID, while lower values indicate OOD cases. In practice, a thresholding method can be applied for effective OOD detection depending on different tasks.
@@ 4.3 Refined Prompt for Near-OOD
Identifying near-OOD cases poses a significant challenge for conventional OOD detection methods, as unseen objects often share closely related semantic meanings with certain in-distribution (ID) categories, making separation difficult. However, off-the-shelf generative and CLIP models inherently possess the ability to distinguish such cases because of their extensive pre-training on large, diverse datasets. To validate this, we proposed a simple approach that, given a predefined set of ID labels, we use knowledge graphs or taxonomies from large language models to retrieve closely related non-ID concepts and explicitly instruct the generative model not to inpaint these concepts as part of the conditioning. For example, if the ID label is “horse”, we can retrieve highly related but OOD classes such as—donkey, zebra, mule, pony, and camel—and explicitly condition the inpainting model to generate a horse that remains distinct from these related categories. We believe this “refined” prompting technique not only supports RONIN to generate outcomes that align better with the ID label but also improve its ability to do near-OOD detection.
@@ 5 Experiments
In-Distribution Method MS-COCO OpenImages FPR@95 $(\downarrow)$ AUROC $(\uparrow)$ FPR@95 $(\downarrow)$ AUROC $(\uparrow)$ PASCAL-VOC ODIN [liang2018odin] 41.65 88.22 55.87 86.46 Energy Score [liu2020energy] 29.48 90.26 24.57 91.24 Mahalanobis [xiao2021we] 63.30 83.26 45.22 87.11 KNN [Sun2022knn] 58.56 85.52 45.00 84.62 VOS [du2022vos] 48.15 88.75 54.63 83.65 SIREN [du2022siren] 64.70 78.68 66.69 75.12 MCM [ming2022delving] 62.47 83.15 71.52 81.45 CLIPN [wang2023clipn] 43.09 85.45 41.74 89.31 TAG [liu2024tag] 61.03 78.35 48.48 87.69 OLE [ding2024zero] 54.23 86.13 49.13 88.07 GL-MCM [miyai2025gl] 78.56 78.64 69.35 82.89 RONIN (5 steps) 27.94 92.35 20.00 92.32 RONIN (10 steps) 29.07 91.10 20.20 92.80 RONIN (15 steps) 25.57 91.63 19.87 92.84 RONIN (20 steps) 25.36 92.31 18.91 93.10 BDD-100k ODIN [liang2018odin] 96.51 55.18 95.56 57.87 Energy Score [liu2020energy] 71.75 74.49 53.33 73.09 Mahalanobis [xiao2021we] 31.75 90.74 23.33 93.81 KNN [Sun2022knn] 35.87 86.58 31.11 92.75 VOS [du2022vos] 65.45 78.34 59.23 80.42 SIREN [du2022siren] 42.86 89.37 37.97 91.78 MCM [ming2022delving] 95.56 55.82 92.22 57.05 CLIPN [wang2023clipn] 28.49 92.14 44.76 85.78 TAG [liu2024tag] 46.43 90.85 55.30 76.67 OLE [ding2024zero] 57.14 78.14 40.00 85.36 GL-MCM [miyai2025gl] 96.83 50.82 94.44 58.88 RONIN (5 steps) 27.94 92.59 26.67 91.23 RONIN (10 steps) 26.67 92.73 25.56 91.33 RONIN (15 steps) 26.03 92.83 24.44 91.65 RONIN (20 steps) 26.03 92.90 23.33 91.90
Figure 4: Side-by-side quantitative visualization. Good cases show synthesized objects consistent with predicted labels and clear OOD score separation; bad cases show inpainting failures or OOD objects too resembling the originals, leading to ineffective OOD score.
@@ 5.1 Experimental Setups
Datasets.   We use Pascal-VOC [everingham2010pascal] and Berkerly DeepDrive (BDD-100k) [yu2020bdd100k] as ID data, and MS-COCO [lin2014microsoft] and OpenImages [kuznetsova2020open] as OOD data. The combinations of these data covered indoor, outdoor, and in-the-wild objects and contexts, allowing us to evaluate RONIN comprehensively. We follow the splits from [du2022siren, du2022vos], where overlapping categories in the ID and OOD datasets are removed. We select a subset of 200 images from the test set of BDD-100k and 400 each from the test sets of Pascal-VOC, MS-COCO, and OpenImages. Data processing details can be found in Appendix A.
@@
Baselines.   We consider three types of baselines: (1) Discriminative zero-shot approaches: MCM [ming2022delving], CLIP-based ODIN [liang2018odin], CLIP-based Energy Score [liu2020energy], CLIPN [wang2023clipn], TAG [liu2024tag], OLE [ding2024zero] and GL-MCM [miyai2025gl]. (2) Alternative generative zero-shot approaches: We first generate synthetic data and then apply standard OOD detection methods such as Mahalanobis [xiao2021we] and KNN [Sun2022knn]. (3) Training-based approaches: for a complete comparison, we also include two representative training-based object-level methods VOS [du2022vos] and SIREN [du2022siren], resulting in a total of eleven baselines. Implementation details can be found in Appendix B
Implementation Details.    For RONIN, we perform class-wise inpainting with Stable Diffusion 2 Inpainting [rombach2021highresolution] with 20 steps. Each object is masked with a center mask covering 0.9 of the height and width of its original bounding box. Object-label similarities are determined by the distance between OpenCLIP [gabriel2021openclip] features, and object-object similarities by the SimCLRv2 [chen2020simclrv2] features.
For the main experiments in Sec. 5.2, we use detected objects obtained from Deformable-DETR [zhu2020deformable], as implemented by [du2022siren], to enable direct comparisons with training-based baselines. However, it is important to note that RONIN is compatible with various types of detectors. In Sec. 5.3, we demonstrate RONIN ’s performance using detections from the pre-trained open-vocabulary detector GroundingDINO [liu2023grounding], showcasing its flexibility and compatibility. The Triplet Similarity (Eq. 4) is computed with $\bm{\alpha}=2$ and $\bm{\beta}=1$. All experiments were conducted on a single NVIDIA GeForce RTX 2080Ti 11GB.
5.2 Quantitative and Qualitative performance
Tab. 1 shows the performance of RONIN and baselines based on DeformableDETR detections across four settings. RONIN achieves the highest results in three scenarios, even with just 5 denoising steps, outperforming training-based methods like VOS and SIREN without retraining on in-domain data. Increasing the number of denoising steps further improves performance, demonstrating that RONIN consistently surpasses all baselines across most settings.
Fig. 4 presents side-by-side visualizations of RONIN with Pascal-VOC as the ID setting. Overall, ID objects are well reconstructed while OOD objects appear distinctly different, enabling effective separation via the triplet score. However, we identify two common types of failures: inpainting failures, where small, occluded, or poorly lit objects prevent meaningful reconstruction, and similarity-score failures, where generated objects remain visually or semantically too close to the original, reducing the effectiveness of the triplet score for OOD detection.
@@ 5.3 Near-OOD Detection
Figure 5: RONIN performance on near-OOD with refined prompting. With distinct inpainting by providing more context, RONIN is able to yield lower scores for near-OOD detection.
To demonstrate the potential of RONIN for near-OOD detection, we constructed a small-scale dataset featuring visually and semantically similar ID and OOD objects. Specifically, we curated 411 samples by pairing ID objects from PascalVOC with closely related OOD counterparts from MS-COCO and OpenImages, forming categories such as “horse vs. zebra”, “dog vs. fox”, and “cat vs. raccoon/tiger”. Using the open-vocabulary detector GroundingDINO [liu2023grounding], we obtained 245 ID and 265 OOD predictions. For the “refined” prompt, we simply leveraged ChatGPT (GPT-4) to generate five nearest non-ID concepts for each ID label as exclusions, with the prompt structure and results are presented in Fig. 5.
Table 2: Near-OOD detection performance. With more information, the “refined” prompt performs better on challenging cases.
FPR@95 ($\downarrow$) AUROC ($\uparrow$) RONIN (Simple prompting) 7.79 96.86 RONIN (Refined prompting) 5.66 98.03
Tab. 2 presents the near-OOD performance under two prompting strategies. Further shown in Fig. 5, the “refined” prompt enables the inpainting model to generate outputs more closely aligned with the predicted labels, thereby improving near-OOD detection. These results support our hypothesis that pre-trained generative and CLIP models already possess the capability for fine-grained OOD detection, given sufficient context.
@@ 6 Ablation Studies
Performance-Speed Trade off.   Fig. 6 analyzes the trade-off between RONIN ’s performance and runtime by varying the number of denoising steps and the image resolutions used during the inpainting process. For relative performance comparison, we include CLIPN [wang2023clipn] as the best-performing baseline, though we do not compare absolute runtimes between the two methods. Reducing the number of denoising steps and downscaling image sizes significantly accelerates RONIN, with only minor performance degradation. These results highlight the flexibility of our approach and its potential suitability for low-resource compute environments, with reasonable performance drops.
@@
Table 3: RONIN with object-wise and class-wise inpainting. Class-wise inpainting retains performance with faster runtime.
@@
FPR@95 $(\downarrow)$ AUROC $(\uparrow)$ Object-wise Class-wise Object-wise Class-wise VOC - COCO 26.60 25.36 91.48 92.31 VOC - OpenImages 20.00 18.91 92.68 93.10 BDD - COCO 28.89 26.03 93.74 92.90 BDD - OpenImages 26.67 23.33 92.42 91.90
Class-wise vs. Object-wise Inpainting.   Tab. 3 compares the “object-wise” inpainting strategy, which processes object individually, with our default “class-wise” approach, which inpaints objects categorically. By leveraging label-wise grouping, class-wise inpainting achieves similar performances while offering more efficient computation cost.
Table 4: Ablation on triplet similarity. All three similarities are essential for effective score (underlined). Emphasizing visual-language semantic alignment ($\alpha$) improves performance (bolded).
[…]
Figure 6: Performance-speed tradeoff of RONIN. Reducing image size and denoising steps speeds up inference with minimal performance loss. At full settings, RONIN achieves top accuracy with competitive runtime. CLIPN is for performance comparison.
Triplet Similarity Measurement.   Tab. 4 presents ablation studies on the triplet similarity score (Eq. 4). The upper half (rows 1–4) assesses the effect of removing individual similarity components, confirming that all are essential for optimal performance. The lower half (rows 5–7) explores the weighting between semantic alignment $\textit{similarity}{(ori,\ \hat{y})}$ (weighted by $\bm{\alpha}$) and visual alignment $\textit{similarity}{(ori,\ inp)}$ (weighted by $\bm{\beta}$). As shown in Fig. 3, emphasizing semantic alignment ($\bm{\alpha}>\bm{\beta}$) yields the best results, guiding our default choice ($\bm{\alpha}=2$, $\bm{\beta}=1$). However, the sensitivities of these weights suggest potential adaptability to more challenging settings.
Table 5: Trade-off performance with different mask sizes. Sufficient mask size (80% bounding box) retains cues for distinguishing object synthesis toward effective OOD performances.
Covered masking ratio $m$ MS-COCO/OpenImages
FPR@95 $(\downarrow)$ AUROC $(\uparrow)$
75% 30.10 / 21.74 90.43 / 92.41
80% 25.80 / 17.74 91.32 / 93.84
100% 29.68 / 20.29 91.61 / 92.61
Impact of Masking Ratios.   Tab. 5 examines the impact of mask size on OOD performance. While smaller masks lead to minimal inpainting changes for effective OOD object synthesis, larger masks do not retain sufficient information for accurate ID object reconstruction, leading to suboptimal performance. More analysis is provided in Sec. C.1.
Table 6: RONIN performances with few-step generative models. The on-par performances demonstrate RONIN effectiveness with minimal reliance on specific models or synthesis quality.
COCO / OpenImages FPR@95 ($\downarrow$) AUROC ($\uparrow$) Stable Diffusion 2 (5 steps) 27.94 / 20.00 92.35 / 92.32 InstaFlow (1 steps) 31.34 / 25.74 88.03 / 90.35
Robust across Generative Models.   Tab. 6 evaluates RONIN across various generative models, focusing on the few-step regime. RONIN shows strong performance even with the one-step generative model InstaFlow [liu2023instaflow], suggesting minimal reliance on specific models or inpainting quality for effective OOD detection. Additional ablations in the many-step regime are provided in Sec. C.2.
Table 7: Recall and falsy reject ratio over ID data, showing RONIN minimal impact on ID objects.
ID set ID object Recall ID Rejected
Pascal-VOC 1007 91.57% 8.43%
BDD 2632 96.67% 4.33%
End-to-end impact on ID object.   Tab. 7 reports Recall and ID rejection ratios after the detection stage, showing that RONIN maintains stable OOD detection without degrading ID accuracy. While BDD mainly contains small, similar vehicle categories, Pascal VOC spans far more diverse objects and animals, making it a more challenging dataset.
7 Conclusion
Overall, RONIN delivers consistently strong performance across diverse settings, from indoor scenes to in-the-wild scenarios, demonstrating clear flexibility and generality. By exploiting subtle misalignments between generative and discriminative outputs, RONIN achieves robust OOD performance without reliance on specific diffusion models or inpainting results, and it remains adaptable to a wide range of object detectors as a training-free, plug-and-play solution. Although current diffusion models limit real-time deployment, many practical applications operate in offline or post-processing pipelines where accuracy and generalizability matter most—areas where RONIN excels. With strong generality and high adaptability across experiments, RONIN offers a meaningful step forward for zero-shot out-of-distribution object detection.
Appendix A Data processing
Following [du2022vos, du2022siren], we use PascalVOC [everingham2010pascal] with 20 categories [^1] PascalVOC ID labels: Person, Car, Bicycle, Boat, Bus, Motorbike, Train, Aeroplane, Chair, Bottle, Dining Table, Potted Plant, TV Monitor, Couch, Bird, Cat, Cow, Dog, Horse, Sheep. and BDD100k [yu2020bdd100k] with 10 categories [^2] BDD100k ID labels: Pedestrian, Rider, Car, Truck, Bus, Train, Motorcycle, Bicycle, Traffic Light, Traffic Sign.. For OOD data, we sample from MS-COCO [lin2014microsoft] and OpenImages [kuznetsova2020open], ensuring no overlap by removing all images containing ID categories. Additionally, to improve alignment with CLIP’s visual-textual representations, we standardize certain British labels (e.g., “Aeroplane”, “Couch”) to their US counterparts (“Airplane”, “Sofa”).
As our method is zero-shot and requires no training, all experiments are performed on subsets drawn from the test sets of VOC, BDD100k, COCO, and OpenImages. To ensure fairness, thoroughness, and a balanced ID–OOD distribution, we sample 200 images from the BDD100k test set and 400 images each from PascalVOC, MS-COCO, and OpenImages. The smaller sample size for BDD100k accounts for its higher object density per image.
Appendix B Implementation Details for Baselines
For discriminative zero-shot baselines, we adapt MCM [ming2022delving], CLIPN [wang2023clipn], TAG [liu2024tag], OLE [ding2024zero], and GL-MCM [miyai2025gl]—originally designed for image-level OOD detection—to the object level by cropping each detection and treating it as an individual image. Each baseline is then applied directly on top of the object detector’s predictions using its default setting, configuration, or conditioning prompt, without any specific modification. ODIN [liang2018odin] and Energy Score [liu2020energy], developed for non-zero-shot classifier-based settings, are adapted similarly by using CLIP for zero-shot classification of detection crops. For generative baselines, we synthesize objects using Stable Diffusion 2, extract features via SimCLRv2 [chen2020simclrv2], and apply standard OOD detectors such as Mahalanobis [xiao2021we] and KNN [Sun2022knn] for fair comparison. For SIREN [du2022siren] and VOS [du2022vos], we follow their original protocols using features from their trained detectors.
Appendix C Additional Ablation Studies
Figure 7: RONIN inpainting performance across masking ratios. Small masks preserve object structure, yielding high similarity for both ID and OOD. Large masks cause excessive distortion, lowering similarity in both cases. A balanced 80–90% ratio retains ID integrity while amplifying OOD deviation, enabling clear ID-OOD separation.
C.1 Impact of Masking Ratios
Extending the ablation study in Tab. 5, we study the impact of the masking ratio on OOD detection by varying the inpainting mask size, covering from 25% to 100% of the predicted bounding box. Quantitative and qualitative results in Tab. 8 and Fig. 7 suggest that despite being robust across masking ratios, RONIN achieves optimal performance when 80–90% of the object is masked for inpainting. As shown in Fig. 7, low masking ratios preserve strong original objects’ appearances in both ID and OOD cases, leading to high similarity scores. In contrast, too large a masking ratio leads to excessive deviation and ambiguity in both cases, leading to poor similarity scores. Covering 75–90% strikes a balance, as the inpainting outcomes not only retain similar inpainting for ID samples but also produce vivid dissimilar inpaintings for OOD ones, enabling RONIN to distinguish ID and OOD objects effectively.
Table 8: Different OOD detection performances when varying mask size for inapinting. RONIN maintains consistent and robust performance across different masking ratios, with the best performance attained when 80%-90% of each object is masked.
Covered masking ratio $m$ MS-COCO OpenImages
FPR@95 $(\downarrow)$ AUROC $(\uparrow)$ FPR@95 $(\downarrow)$ AUROC $(\uparrow)$
25% 42.47 88.53 36.74 90.40
50% 34.23 89.39 23.91 92.45
75% 30.10 90.43 21.74 92.41
80% 25.80 91.32 17.74 93.84
90% 25.57 91.53 19.57 92.46
100% 29.68 91.61 20.29 92.61